# Leveraging Pre-Trained Visual Transformers for Multi-Band Photometric Light Curve Classification

### v2

@@ 2.2 Visual transformer models
SwinV2 also replaces the relative position bias with a log-spaced continuous position bias. This approach is particularly relevant in hierarchical models like SwinV2, where the patch merging operation, detailed in the following paragraph, changes the token resolution at each stage of the network. The log-spaced continuous position bias provides a flexible encoding of spatial relationships that adapts naturally as the resolution changes, allowing the model to maintain meaningful attention patterns at multiple scales. In practice, this helps the model assign more attention to neighboring regions (where fine details or local patterns often reside) while remaining robust to global variations in the visual structure of different light curves.
@@ 3.1.1 MACHO labeled
The public Catalog of Variable Stars [^2] MACHO Labeled Data of the Variable Stars. (Alcock et al. 2003) contains labeled light curves from the MACHO survey, which observed variable stars in the Large Magellanic Cloud (LMC) and the Galactic Bulge over multiple years, capturing their photometric variability in the B and R bands. These light curves are represented in magnitude space. The catalog labels were assigned through automated classification procedures based on statistical properties of the light curve morphology, such as period, amplitude, and Fourier decomposition coefficients. While visual verification was commonly employed in some of the MACHO team’s publications (Alcock et al. 1996, 1997a, 1997b, 1999, 2002), there is no evidence in the literature confirming that a systematic visual inspection was performed for the full catalog used in this work. Therefore, we assume that the classification relied solely on the extracted features. It is important to note that we do not use any auxiliary features or pre-computed variables at any stage of the preprocessing or training pipeline, which could otherwise lead to a trivial classification task. This dataset also has a median cadence of $0.75$ days with a standard deviation of $5.03$ days, considering both bands. The labeled portion we use corresponds to the same one used in Donoso-Oliva et al. (2023), which was categorized into six distinct classes, comprising 20 89420\,894 variable stars. We will refer to this as the Full dataset.
Following the approach in Donoso-Oliva et al. (2023), we use scenarios with few samples per class (spc) to train the model and assess its ability to learn from limited labeled data while keeping the test set unchanged. From the Full dataset, we extracted 20 spc and 500 spc as our new training datasets. As in the referenced study, the IDs for each fold were selected randomly from the Full dataset. For example, the IDs for 20 and 500 spc in the first fold differ from those in the second fold. Our training, validation, and test sets in each fold use the same IDs that were used in the referenced study. The test set contains $100$ light curves per class and remains fixed across all scenarios (20 spc, 500 spc, and Full dataset). The training dataset is divided into three stratified folds, maintaining an 80/20 split for training and validation. This results in $96$ and $24$ light curves for the 20 spc scenario, 2 4002\,400 and $600$ for the 500 spc scenario, and 16 23216\,232 and 4 0624\,062 for the Full dataset. Figure 3 illustrates the class distribution in the training and validation sets for the Full dataset.
@@
3.1.2 ELAsTiCC Challenge
The second dataset used in this study comes from the ELAsTiCC [^3] ELAsTiCC Challenge Data Portal. challenge (Narayan & Team 2023), designed to simulate the observations of the future LSST at the Vera C. Rubin Observatory. It contains 1 845 1461\,845\,146 simulated light curves, each represented across six bands: u, g, r, i, z, and Y, covering the visible and near-infrared spectrum. The data includes multiple observations of each object spread over several simulated years. ELAsTiCC spans a diverse array of transient and variable stellar events, categorized into $32$ distinct classes, with a median cadence of $0.93$ days and a standard deviation of $16.18$ days, considering all six bands. Specifically, we used the first ELAsTiCC campaign, which was used in Cabrera-Vives et al. (2024), and adopted the same class distribution as their study, encompassing $20$ different categories. Additionally, similar to the approach outlined in the referenced study, we used the light curves in flux space and applied the PHOTFLAG [^4] Description of the data variables from the first round of ELAsTiCC. key to filter out saturated data, which also helps exclude spurious detections outside the main event, particularly for transient objects. For clarity, we note that no derived features from the light curves or metadata were used at any stage of the training process. For each light curve, we also considered both alerts and forced photometry points, beginning $30$ days prior to the first alert up to the final detection. Non-detections following the last detection were omitted from the dataset.
@@ 3.2 Implementation details
The conversion of light curves to images was performed using the Matplotlib Python package, and the images were directly transformed into tensors with RGB channels. We generated images with dimensions of $256{\times}256$, as this is the size for which the VT model was pre-trained. In both the Grid and Overlay approaches, the images were adjusted to minimize extra space, and in the Grid approach, borders were added to separate the bands (see Fig. 1 for an illustration of these representations).
@@
We experimented with varying key hyperparameters during the conversion of light curves into images to identify the best visual representation that enhances classification performance, as measured by the F1-score, which balances how accurately the model identifies light curves of a given class (precision) and its ability to recognize all light curves that truly belong to that class (recall). Table 1 summarizes the hyperparameters used. Since the MACHO labeled dataset contains a limited number of light curves and to assess the model’s ability to learn from datasets with small to medium sample sizes per class, we performed hyperparameter tuning using grid search and k-fold cross-validation for each scenario (20 spc, 500 spc, and Full dataset). This means that for each fold, a total of $160$ combinations were tested for the 20 spc, 500 spc, and Full dataset. In contrast, due to the considerable size of the ELAsTiCC dataset, we opted to execute hyperparameter tuning with just one validation fold. This decision stemmed from the high computational expense associated with evaluating multiple folds and the anticipated low variance across folds in large-scale datasets. As before, a total of $160$ combinations were tested, but only in one fold. The lower and upper bounds for the hyperparameters marker size and line width were determined based on the degree of visual overlap between the observations and lines in the images. Figure 1 (a) and (b) present images generated using the best hyperparameters identified for the Grid and Overlay approaches, respectively.
@@ 4 Results
Table 2: F1-score (%) on the MACHO test set for models trained with 20 and 500 samples per class (spc), and the Full dataset, comprising 20 29420\,294 labeled light curves used for training.
@@
 [^7] F denotes a frozen transformer, where only the classifier is updated. T represents a trainable transformer, where both the transformer and classifier are updated during training.
@@ 4.1 Model performance in single- and multi-band light curves
To evaluate the classification performance of our approach using a single band, we compared SwinV2 with Astromer. Our initial experiments focused on the R band of the MACHO dataset to ensure direct comparability with the results reported in the referenced study. We conducted hyperparameter tuning to optimize the model’s performance and to assess its ability to learn from datasets with small to medium sample sizes per class. The best combination of hyperparameters, using one band, for the 20 and 500 spc used to train the model, as well as the Full dataset, can be found at the top of Table 5 in Appendix B. These combinations were then used to generate the SwinV2 results presented in Table 2.
Figure 5: SwinV2 classification performance on the MACHO test set using single- and multi-band data, trained on varying sample sizes. The figure presents the best results from the hyperparameter search for both the Overlay and Grid approaches.
Table 2 presents the single-band classification results on the MACHO dataset in the R band for SwinV2 and Astromer. The table reports the F1-scores along with their standard deviations for SwinV2 (in trainable configurations) and Astromer (in both frozen and trainable configurations) across different dataset sizes. In this context, a trainable model refers to one in which both the transformer and classifier layers are updated during training, whereas a frozen model retains the transformer’s pre-trained weights and only updates the classifier. Since results for the Full dataset were not available for Astromer, we compared it only in the limited-data scenarios (20 and 500 spc). However, the Full dataset results provide a basis for evaluating our multi-band method. As shown in Table 2, SwinV2 consistently outperformed Astromer across all configurations and dataset sizes. For 20 spc, SwinV2 achieved an F1-score of $60.1\%$, which is $4.1\%$ points higher than Astromer’s best result of $56.0\%$. For 500 spc, SwinV2 attained $77.4\%$, surpassing Astromer by $4.4\%$ points. These results demonstrate the robustness and adaptability of the VT model in achieving high classification performance on light curves, particularly in low-data scenarios. When tested on the unbalanced dataset (i.e., the Full dataset, as illustrated in Fig. 3), SwinV2 achieved an F1-score of $79.8\%$, surpassing its performance on the balanced dataset with 500 spc ($77.4\%$). This result highlights the VT model’s ability to maintain strong performance even in the presence of data imbalance.
@@
To demonstrate the impact of the multi-band approach, we extended our analysis to incorporate both the R and B bands from the MACHO dataset, employing two distinct plotting strategies: Grid and Overlay. As before, we performed hyperparameter tuning on each scenario. The best combination of hyperparameters using two bands can be seen in the middle of Table 5 in Appendix B. The results in Fig. 5 show that incorporating multi-band inputs improves F1-scores compared to single-band experiments. Using two bands led to performance improvements of $6.8\%$, $4.3\%$, and $4.5\%$ for the 20, 500, and Full dataset settings, respectively, relative to the best multiband approach in each case. These improvements highlight the model’s ability to effectively leverage the additional information provided by multiple bands, reinforcing the value of multi-band configurations for capturing the complex behavior of light curves through image-based representations. For 20 spc, Overlay outperformed Grid (F1-score: $66.9\%$ vs. $64.3\%$). With 500 spc, Grid achieved an F1-score of $81.7\%$, marginally surpassing Overlay ($80.6\%$). When using the Full dataset, Grid attained a higher F1-score of $84.3\%$, slightly outperforming Overlay ($82.3\%$). These results indicate that Grid achieved better performance in two configurations. However, the differences were minimal.
@@
The confusion matrices for the best-performing SwinV2 models across different dataset sizes, for both single- and multi-band settings, are shown on the left and right of Fig. 6, respectively. These matrices illustrate the recall per class, providing insight into how well the model correctly identifies each category. For the subset with 20 spc, using two bands instead of one improved recall in 4 out of 6 classes (RRab, Cep_1, EC, and LPV) but decreased it for RRc and Cep_0. For the subset with 500 spc, recall increased in 5 out of 6 classes (RRab, Cep_0, Cep_1, EC, and LPV) but decreased by $7\%$ for RRc. For the Full dataset, a similar pattern was observed. Recall improved in 4 out of 6 classes (RRc, Cep_0, Cep_1, and LPV), remained unchanged for EC, and decreased by $1\%$ for RRab. The consistent improvement observed in Cep_1, EC, and LPV suggests that these classes benefit from the additional spectral information. In contrast, RRab, RRc, and Cep_0 exhibit mixed results, indicating that while multi-band inputs are beneficial in most cases, they do not always provide a clear advantage, especially for certain classes that remain challenging to distinguish due to their similar flux variations in both bands. This is particularly evident in RRc and RRab, as well as Cep_0 and Cep_1.
@@
Figure 7 shows the distribution of F1-scores obtained during the hyperparameter tuning process for the MACHO dataset, considering both single-band and two-band approaches for each scenario (20 spc, 500 spc, and the Full dataset). In general, the median F1-score values across all combinations remain consistent with those of the best combinations presented in Fig. 5, where the Overlay approach performed better for 20 spc, while the Grid approach was superior for 500 spc and the Full dataset. In particular, the F1-score for 500 spc and the Full dataset exhibited little variation, regardless of the hyperparameter combinations used. The exception is the 20 spc scenario, where a poor choice of hyperparameters can lead to low performance due to the limited amount of data, which restricts the model’s ability to extract sufficient discriminative information from the light curves.
@@
Figure 7: Distribution of F1-scores obtained during hyperparameter tuning for the MACHO dataset, comparing single-band and two-band (Overlay and Grid) approaches across different dataset sizes (20 spc, 500 spc, and the Full dataset).
@@
To explore this aspect, we conducted a complementary experiment using a brightness-biased sampling strategy. Specifically, we modified only the training set by selecting the brightest objects of each class from the Full dataset, based on their median magnitudes, while maintaining the 20 and 500 samples per class configurations. We refer to these settings as 20 spc-b and 500 spc-b, where the suffix -b denotes brightness-biased sampling. The test set remained unchanged from previous experiments. This introduces a distribution shift between training and testing, as the training set includes only the brightest sources, while the test set covers the full magnitude range. The results, presented in Table 3, show lower F1-scores compared to those obtained with randomly sampled training data. For 20 spc-b, the F1-score dropped from 60.1% to 48.9% in the one-band configuration, from $66.9\%$ to $55.8\%$ in the Overlay approach, and from $64.3\%$ to $51.6\%$ in the Grid approach. For 500 spc-b, the F1-score decreased from $77.4\%$ to $69.5\%$ in one-band, from $80.6\%$ to $72.9\%$ in Overlay, and from $81.7\%$ to $74.7\%$ in Grid. Despite these differences, the model still achieves relatively high scores, particularly with two-band inputs. Among the visual strategies, Overlay performs better under smaller training sizes, while Grid yields the highest performance with 500 spc-b. These findings indicate that the model remains effective even under sampling conditions that reflect realistic observational biases.
4.2 Model performance in ELAsTiCC Challenge
Table 4: F1-score (%) on the ELAsTiCC test set for ATAT using only light curves and for SwinV2 with the Overlay and Grid approaches.

### v4

@@ 2.2 Vision Transformer models
SwinV2 also replaces the relative position bias with a log-spaced continuous position bias. This approach is particularly relevant in hierarchical models such as SwinV2, where the patch merging operation, detailed in the following paragraph, changes the token resolution at each stage of the network. The log-spaced continuous position bias provides a flexible encoding of spatial relationships that adapts naturally as the resolution changes, allowing the model to maintain meaningful attention patterns at multiple scales. In practice, this helps the model assign more attention to neighboring regions (where fine details or local patterns often reside) while remaining robust to global variations in the visual structure of different light curves.
@@ 3.1.1 MACHO labeled
The public Catalog of Variable Stars [^2] MACHO Labeled Data of the Variable Stars. (Alcock et al. 2003) contains labeled light curves from the MACHO survey, which observed variable stars in the Large Magellanic Cloud (LMC) and the Galactic bulge over multiple years, capturing their photometric variability in the B and R bands. These light curves are represented in magnitude space. The catalog labels were assigned through automated classification procedures based on statistical properties of the light curve morphology, such as period, amplitude, and Fourier decomposition coefficients. While visual verification was commonly employed in some of the MACHO team’s publications (Alcock et al. 1996, 1997a, 1997b, 1999, 2002), there is no evidence in the literature confirming that a systematic visual inspection was performed for the full catalog used in this work. Therefore, we assume that the classification relied solely on the extracted features. It is important to note that we do not use any auxiliary features or pre-computed variables at any stage of the preprocessing or training pipeline, which could otherwise lead to a trivial classification task. This dataset also has a median cadence of $0.75$ days with a standard deviation of $5.03$ days, considering both bands. The labeled portion we use corresponds to the same one used in Donoso-Oliva et al. (2023), which was categorized into six distinct classes, comprising 20 89420\,894 variable stars. We refer to this as the “full dataset”.
Following the approach in Donoso-Oliva et al. (2023), we used scenarios with a few samples per class (spc) to train the model and assess its ability to learn from limited labeled data while keeping the test set unchanged. From the “full dataset”, we extracted 20 spc and 500 spc as our new training datasets. As in the referenced study, the IDs for each fold were selected randomly from the “full dataset”. For example, the IDs for 20 and 500 spc in the first fold differ from those in the second fold. Our training, validation, and test sets in each fold use the same IDs that were used in the referenced study. The test set contains $100$ light curves per class and remains fixed across all scenarios (20 spc, 500 spc, and “full dataset”). The training dataset is divided into three stratified folds, maintaining an 80/20 split for training and validation. This results in $96$ and $24$ light curves for the 20 spc scenario, 2 4002\,400 and $600$ for the 500 spc scenario, and 16 23216\,232 and 4 0624\,062 for the “full dataset”. Figure 3 illustrates the class distribution in the training and validation sets for the “full dataset”.
@@
3.1.2 ELAsTiCC
The second dataset used in this study comes from the ELAsTiCC [^3] ELAsTiCC Challenge Data Portal. challenge (Narayan & Team 2023), which is designed to simulate the observations of the future LSST at the Vera C. Rubin Observatory. It contains 1 845 1461\,845\,146 simulated light curves, and each is represented across six bands (u, g, r, i, z, and Y) covering the visible and near-infrared spectrum. The data includes multiple observations of each object spread over several simulated years. ELAsTiCC spans a diverse array of transient and variable stellar events, categorized into $32$ distinct classes, with a median cadence of $0.93$ days and a standard deviation of $16.18$ days, considering all six bands. Specifically, we used the first ELAsTiCC campaign, which was used in Cabrera-Vives et al. (2024), and adopted the same class distribution as their study, encompassing $20$ different categories. Additionally, similar to the approach outlined in the referenced study, we used the light curves in flux space and applied the PHOTFLAG [^4] Description of the data variables from the first round of ELAsTiCC. key to filter out saturated data, which also helps exclude spurious detections outside the main event, particularly for transient objects. For clarity, we note that no derived features from the light curves or metadata were used at any stage of the training process. For each light curve, we also considered both alerts and forced photometry points, beginning $30$ days prior to the first alert up to the final detection. Non-detections following the last detection were omitted from the dataset.
@@ 3.2 Implementation details
The conversion of light curves to images was performed using the Matplotlib Python package, and the images were directly transformed into tensors with RGB channels. We generated images with dimensions of $256{\times}256$, as this is the size for which the VT model was pre-trained. In both the “Grid” and “Overlay” approaches, the images were adjusted to minimize extra space, and in the “Grid” approach, borders were added to separate the bands (see Fig. 1 for an illustration of these representations).
@@
We experimented with varying key hyperparameters during the conversion of light curves into images to identify the best visual representation that enhances classification performance, as measured by the F1-score, which balances how accurately the model identifies light curves of a given class (precision) and its ability to recognize all light curves that truly belong to that class (recall). Table 1 summarizes the hyperparameters used. Since the MACHO labeled dataset contains a limited number of light curves and to assess the model’s ability to learn from datasets with small to medium sample sizes per class, we performed hyperparameter tuning using grid search and k-fold cross-validation for each scenario (20 spc, 500 spc, and “full dataset”). This means that for each fold, a total of $160$ combinations were tested for the 20 spc, 500 spc, and “full dataset”. In contrast, due to the considerable size of the ELAsTiCC dataset, we opted to execute hyperparameter tuning with just one validation fold. This decision stemmed from the high computational expense associated with evaluating multiple folds and the anticipated low variance across folds in large-scale datasets. As before, a total of $160$ combinations were tested, but only in one fold. The lower and upper bounds for the hyperparameters marker size and line width were determined based on the degree of visual overlap between the observations and lines in the images. Figure 1 (a) and (b) present images generated using the best hyperparameters identified for the “Grid” and “Overlay” approaches, respectively.
@@ 4 Results
Table 2: F1-score (%) on the MACHO test set for models trained with 20 and 500 samples per class (spc), and the “full dataset” comprising 20 29420\,294 labeled light curves used for training. [^7] F denotes a frozen transformer, where only the classifier is updated. T represents a trainable transformer, where both the transformer and classifier are updated during training.
@@ 4.1 Model performance in single- and multi-band light curves
To evaluate the classification performance of our approach using a single band, we compared SwinV2 with Astromer. Our initial experiments focused on the R band of the MACHO dataset to ensure direct comparability with the results reported in the referenced study. We conducted hyperparameter tuning to optimize the model’s performance and to assess its ability to learn from datasets with small to medium sample sizes per class. The best combination of hyperparameters, using one band, for the 20 and 500 spc used to train the model, as well as the “full dataset”, can be found at the top of Table 5 in Appendix B. These combinations were then used to generate the SwinV2 results presented in Table 2.
Figure 5: SwinV2 classification performance on the MACHO test set using single- and multi-band data trained on varying sample sizes. The figure presents the best results from the hyperparameter search for both the “Overlay” and “Grid” approaches.
Table 2 presents the single-band classification results on the MACHO dataset in the R band for SwinV2 and Astromer. The table reports the F1-scores along with their standard deviations for SwinV2 (in trainable configurations) and Astromer (in both frozen and trainable configurations) across different dataset sizes. In this context, a trainable model refers to one in which both the transformer and classifier layers are updated during training, whereas a frozen model retains the transformer’s pre-trained weights and only updates the classifier. Since results for the “full dataset” were not available for Astromer, we compared it only in the limited-data scenarios (20 and 500 spc). However, the “full dataset” results provide a basis for evaluating our multi-band method. As shown in Table 2, SwinV2 consistently outperformed Astromer across all configurations and dataset sizes. For 20 spc, SwinV2 achieved an F1-score of $60.1\%$, which is $4.1\%$ points higher than Astromer’s best result of $56.0\%$. For 500 spc, SwinV2 attained $77.4\%$, surpassing Astromer by $4.4\%$ points. These results demonstrate the robustness and adaptability of the VT model in achieving high classification performance on light curves, particularly in low-data scenarios. When tested on the unbalanced dataset (i.e., the “full dataset”, as illustrated in Fig. 3), SwinV2 achieved an F1-score of $79.8\%$, surpassing its performance on the balanced dataset with 500 spc ($77.4\%$). This result highlights the VT model’s ability to maintain strong performance even in the presence of data imbalance.
@@
To demonstrate the impact of the multi-band approach, we extended our analysis to incorporate both the R and B bands from the MACHO dataset, employing two distinct plotting strategies: “Grid” and “Overlay”. As before, we performed hyperparameter tuning on each scenario. The best combination of hyperparameters using two bands can be seen in the middle of Table 5 in Appendix B. The results in Fig. 5 show that incorporating multi-band inputs improves F1-scores compared to single-band experiments. Using two bands led to performance improvements of $6.8\%$, $4.3\%$, and $4.5\%$ for the 20, 500, and “full dataset” settings, respectively, relative to the best multi-band approach in each case. These improvements highlight the model’s ability to effectively leverage the additional information provided by multiple bands, reinforcing the value of multi-band configurations for capturing the complex behavior of light curves through image-based representations. For 20 spc, “Overlay” outperformed “Grid” (F1-score: $66.9\%$ vs. $64.3\%$). With 500 spc, “Grid” achieved an F1-score of $81.7\%$, marginally surpassing “Overlay” ($80.6\%$). When using the “full dataset”, “Grid” attained a higher F1-score of $84.3\%$, slightly outperforming “Overlay” ($82.3\%$). These results indicate that “Grid” achieved better performance in two configurations. However, the differences were minimal.
@@
The confusion matrices for the best-performing SwinV2 models across different dataset sizes, for both single- and multi-band settings, are shown on the left and right of Fig. 6, respectively. These matrices illustrate the recall per class, providing insight into how well the model correctly identifies each category. For the subset with 20 spc, using two bands instead of one improved recall in 4 out of 6 classes (RRab, Cep_1, EC, and LPV) but decreased it for RRc and Cep_0. For the subset with 500 spc, recall increased in 5 out of 6 classes (RRab, Cep_0, Cep_1, EC, and LPV) but decreased by $7\%$ for RRc. For the “full dataset”, a similar pattern was observed. Recall improved in 4 out of 6 classes (RRc, Cep_0, Cep_1, and LPV), remained unchanged for EC, and decreased by $1\%$ for RRab. The consistent improvement observed in Cep_1, EC, and LPV suggests that these classes benefit from the additional spectral information. In contrast, RRab, RRc, and Cep_0 exhibit mixed results, indicating that while multi-band inputs are beneficial in most cases, they do not always provide a clear advantage, especially for certain classes that remain challenging to distinguish due to their similar flux variations in both bands. This is particularly evident in RRc and RRab, as well as Cep_0 and Cep_1.
@@
Figure 7 shows the distribution of F1-scores obtained during the hyperparameter tuning process for the MACHO dataset, considering both single-band and two-band approaches for each scenario (20 spc, 500 spc, and the “full dataset”). In general, the median F1-score values across all combinations remain consistent with those of the best combinations presented in Fig. 5, where the “Overlay” approach performed better for 20 spc, while the “Grid” approach was superior for 500 spc and the “full dataset”. In particular, the F1-score for 500 spc and the “full dataset” exhibited little variation, regardless of the hyperparameter combinations used. The exception is the 20 spc scenario, where a poor choice of hyperparameters can lead to low performance due to the limited amount of data, which restricts the model’s ability to extract sufficient discriminative information from the light curves.
@@
Figure 7: Distribution of F1-scores obtained during hyperparameter tuning for the MACHO dataset comparing single-band and two-band (“Overlay” and “Grid”) approaches across different dataset sizes (20 spc, 500 spc, and the “full dataset”).
@@
To explore this aspect, we conducted a complementary experiment using a brightness-biased sampling strategy. Specifically, we modified only the training set by selecting the brightest objects of each class from the “full dataset”, based on their median magnitudes, while maintaining the 20 and 500 samples per class configurations. We refer to these settings as 20 spc-b and 500 spc-b, where the suffix -b denotes brightness-biased sampling. The test set remained unchanged from previous experiments. This introduces a distribution shift between training and testing, as the training set includes only the brightest sources, while the test set covers the full magnitude range. The results, presented in Table 3, show lower F1-scores compared to those obtained with randomly sampled training data. For 20 spc-b, the F1-score dropped from 60.1% to 48.9% in the one-band configuration, from $66.9\%$ to $55.8\%$ in the “Overlay” approach, and from $64.3\%$ to $51.6\%$ in the “Grid” approach. For 500 spc-b, the F1-score decreased from $77.4\%$ to $69.5\%$ in one-band, from $80.6\%$ to $72.9\%$ in “Overlay”, and from $81.7\%$ to $74.7\%$ in “Grid”. Despite these differences, the model still achieves relatively high scores, particularly with two-band inputs. Among the visual strategies, “Overlay” performs better under smaller training sizes, while “Grid” yields the highest performance with 500 spc-b. These findings indicate that the model remains effective even under sampling conditions that reflect realistic observational biases.
4.2 Model performance in ELAsTiCC
Table 4: F1-score (%) on the ELAsTiCC test set for ATAT using only light curves and for SwinV2 with the “Overlay” and “Grid” approaches.